# Gaussian process regression + deep neural network autoencoder for probabilistic surrogate modeling in nonlinear mechanics of solids

### v1

@@ Abstract
Many real-world applications demand accurate and fast predictions, as well as reliable uncertainty estimates. However, quantifying uncertainty on high-dimensional predictions is still a severely under-invested problem, especially when input-output relationships are non-linear. To handle this problem, the present work introduces an innovative approach that combines autoencoder deep neural networks with the probabilistic regression capabilities of Gaussian processes. The autoencoder provides a low-dimensional representation of the solution space, while the Gaussian process is a Bayesian method that provides a probabilistic mapping between the low-dimensional inputs and outputs. We validate the proposed framework for its application to surrogate modeling of non-linear finite element simulations. Our findings highlight that the proposed framework is computationally efficient as well as accurate in predicting non-linear deformations of solid bodies subjected to external forces, all the while providing insightful uncertainty assessments.
@@ 1 Introduction
Rapid advancements in machine learning has naturally fueled developments to approximate modelling. There has been a widespread adoption of machine learning techniques to accelerate computationally intensive numerical methods [11, 12, 13, 14, 15]. In particular, deep learning (DL) models have been successfully used as accurate and computationally inexpensive surrogates to solve non-linear problems in mechanics [16, 17, 18, 19, 20]. In context of ROM, DL methods are being extensively used to find non-linear reduced order representations of the high dimensional data [21]. At the same time, they are computationally inexpensive at the prediction stage, hence, they are highly suitable in a wide area of applications requiring real time simulations [22, 23].
@@
The framework proposed in this contribution focuses on alleviating part of the scaling issue by directly reducing the dimension of outputs, $M$. This is achieved by applying the compressed latent representation of outputs via, so called, autoencoder deep neural networks—a strategy akin to approaches found in related works [43, 44]. Autoencoder neural networks, as an unsupervised learning technique, have been commonly adopted to efficiently learning compressed representations across various domains [45, 46, 47], also in the context of computational fluid dynamics [48, 49] and computational solid mechanics [50, 51]. The primary advantage of autoencoder DNNs is their ability to reduce the dimensionality of data in a non-linear way [52]. As such, they perform much better than POD or PCA when high non-linearities are present in data [53], and are thus used in the present work as an integral component of the proposed probabilistic framework.
In summary, in this work we combine GPR with autoencoder DNNs to provide surrogate model for force-displacement predictions in mechanics of solids. In the proposed framework, we reduce the high-dimensional 2D/3D finite element solutions (full-field displacement) into their reduced representation by applying autoencoder neural networks, and then use GPR to provide probabilistic mapping between the input forces and the reduced-order displacements. During the prediction phase, for a given unseen force, GPR predicts the reduced-order (latent space) displacement which is then projected to the full field using the decoder part of autoencoder network. We demonstrate the robustness and versatility of our framework by applying it to synthetic datasets generated from the non-linear finite element simulations.
@@ 2.1 Obtaining latent representations with autoencoder neural networks
${\bf u}^{l}=\mathcal{U}_{\text{encoder}}({\bf u},\bm{\theta_{\text{auto}}}^{*})$ (5)
@@ 2.2 Gaussian Process Regression in the latent space
As discussed in the introduction, GPs face scalability challenges as the number of their inputs (data points), $N$, and dimension of outputs, $M$, increase. In this work we focus on reducing the bottleneck related to $M$. It is common practice to implement independent GPs for each dimension of the output, assuming that all output components are independent of each other. This approach simplifies the complexity of a multi-output GPR problem from $\mathcal{O}(N^{3}M^{3})$ to $M$ problems with $\mathcal{O}(N^{3})$ complexity (i.e., $\mathcal{O}(N^{3}M)$), where $N$ is the number of data points and $M$ is the dimensionality of the output. However, for problems with a large number of degrees of freedom (DOF), such as the examples in this contribution, even $\mathcal{O}(N^{3}M)$ scaling proves to be computationally costly since $M$ independent GPRs must be implemented for each dimension of the outputs. Therefore, in this study, we project the high-dimensional displacements (with dimension $M$) to a significantly lower-dimensional space ($L\ll M$) and then employ $L$ independent GPs to establish a probabilistic relationship between input forces and the latent representation of displacements, which scales as $\mathcal{O}(N^{3}L)$. Because our framework uses independent GPRs for each latent output, we provide a brief and practical introduction to GPs in this subsection. Interested readers are referred to [29] for more details.
@@
Note that, $p(\mbox{\boldmath{$y$}}|{\bf F},\bm{\theta},\sigma^{2})$ in Eq. (12), is the conditional likelihood of observations given the parameters set $\bm{\theta_{\text{GP}}}$, the input matrix ${\bf F}$, $\sigma^{2}$, hence one also can use Bayesian inference for identification of parameters by setting priors on unknowns (i.e., $\bm{\theta_{\text{GP}}}$, and $\sigma^{2}$, see [56]). Moreover, GP and autoencoder network are trained separately. Autoencoder network needs to be trained first in order to generate the latent state dataset, which is then used to train the GP.
@@ 2.3 Projecting latent GP predictions to the full field space using decoder
As depicted in the Figure 2, latent space predictions obtained using GP are projected to full space using decoder part of the autoencoder network. At the inference stage, for a new input force ${\bf f^{*}}$, GP first generates latent displacement distribution $p({\tilde{u}}^{*l}|{\bf f^{*}})$ for each latent output, which is a Gaussian distribution with the mean and variance obtained using Eq.(10) and Eq.(11) respectively.
$p({\tilde{u}}^{*l}|{\bf f^{*}})=\mathcal{N}({\bf\mathbb{E}}_{\text{GP}},{\bf\mathbb{V}}_{\text{GP}}),$ (13)
Subsequently, the distributions of full field displacements are obtained by translating samples from the latent predictive distributions through the decoder component. Let ${\bf u}_{s}^{l}$ denote the vector of samples drawn from distributions corresponding to all latent outputs, obtained using independent GPs for an input ${\bf f}^{*}$. The displacement corresponding to the full field solution for each degree of freedom (DOF) is a Gaussian distribution, whose mean and variance are computed as follows:
$\begin{split}(\tilde{u}^{*}_{\mu})_{i}&\approx\frac{1}{S}\sum_{s=1}^{S}(\mathcal{U}_{\text{decoder}}({\bf u}_{s}^{l}))_{i}\\ {(u_{\sigma}^{*})_{i}}^{2}&\approx\frac{1}{S-1}\sum_{s=1}^{S}((\mathcal{U}_{\text{decoder}}({\bf u}_{s}^{l}))_{i}-(u^{*}_{\mu})_{i})^{2}\end{split}$ (14)
@@
2.4 Understanding of errors
As depicted in the Figure 2, for a new unseen case, the GP is first used to make displacement prediction in the latent space, which is subsequently projected to full space solution using the decoder part of the autoencoder network. Hence the error in the framework arises from GP fitting as well as the compression and reconstructions steps using the autoencoder network. For the sake of simplicity, in Figure 3, we demonstrate the schematic for a single DOF prediction.
The mean prediction of the framework is considered as the displacement solution obtained using the framework. Hence the absolute difference between this mean prediction and the true FEM solution is considered as the error of the framework, $e_{\text{f}}$, as follows:
$e_{\text{f}}=|\tilde{u}_{\mu}-u|$ (15)
The error of reconstruction is computed by taking absolute difference between the reconstructed solution (represented by) and the FEM solution. Reconstructed solution is obtained by projecting the true latent solution (as obtained by Eq. (5)) to the full field, thus neglecting any contribution of the GP. The reconstruction solution is computed as ${\bf u}_{r}=\mathcal{U}_{\text{decoder}}({\bf u}^{l})$, and the corresponding reconstruction error is (for a particular DOF in this case):
$e_{r}=|u_{r}-u|$ (16)
Finally, the contribution of GP fitting error corresponding to the full field representation is computed by taking the difference between the mean prediction and the reconstruction solution as follows:
$e_{\text{GP}}=|\tilde{u}_{\mu}-u_{r}|$ (17)
Note: All the errors are computed in the decoded space.
@@ 2.5 Finite element formulation for non-linear deformations of solid bodies
Remark: The vector ${\bf f}$ provides a compressed information about the external loading. The solid body can be subjected to either loads applied to a single node, set of nodes, or body forces acting throughout the entire region. When dealing with point loads, we describe the force using its magnitude ($f_{x}$, $f_{y}$) and the position at which the force is applied (relative to the fixed boundary). In scenarios involving body forces, we provide the force density components ($f_{x}$, $f_{y}$) or ($f_{x}$, $f_{y}$, $f_{z}$), for 2D or 3D problems, respectively. See also more details in Section 3.1.
@@ 3.1 Dataset generation information
In this research, we examine two distinct scenarios, as illustrated in Figure 4: a 2D beam and a 3D liver. These scenarios are subjected to varied loading conditions and are simulated using a non-linear finite element method during the data generation phase.
2D beam case: In this scenario, the 2D beam is applied with point forces with random direction and magnitude. These forces are applied to the nodes located along the top line, which is visually represented by red line in the Figure 4a. For the 2D beam case, displacement data is provided for all degrees of freedom (DOFs) of the mesh, capturing the full-field displacements of the beam. However, the forces used as inputs to the surrogate framework are represented as 3-dimensional arrays. Specifically, the applied point load is characterized by two components, namely $f_{x}$ and $f_{y}$, representing the force magnitudes in the x and y directions, respectively. Additionally, the distance, $d$, from the fixed boundary to the point of application is considered. This information is visually depicted in Figure 4a.
3D liver case: In this case, the 3D liver model is subjected to body forces with random directions and magnitudes distributed throughout its volume. The liver is also constrained at specific nodes located at the right end. This simulation is aimed at understanding the deformations of the liver under the influence of these distributed body forces and boundary constraints. Similar to the beam case, displacement data is provided for all DOFs of the liver mesh, representing the full-field displacements of the organ. However, in contrast to the beam case, the forces used as inputs to the surrogate framework are the body force density components in the x, y, and z directions, respectively ($b_{x},b_{y},b_{z}$).
Further details, such as the range of external forces applied, as well as other information used to generate the datasets are given in the Table 1.
@@
Figure 4: Schematics of examples considered in this work (a) 2D beam discretised with quad elements is applied with point loads on the nodes lying on red line, this example has been taken from [54]. (b) 3D liver is applied with random body forces within a prescribed region.
@@ 3.2.1 Autoencoder networks: Architectures & training details
In this study, we employ two types of autoencoder networks. Firstly, we utilize a CNN-based architecture, illustrated in Figure 5, tailored for the 2D beam case. Secondly, we utilize a fully connected autoencoder network, depicted in Figure 6, implemented for the liver case.
Both of these networks consist of an encoder and a decoder component. The primary function of these autoencoder networks is to compress the displacement solutions of the full field space into corresponding latent states. It must be noted that while the autoencoders handle displacement solutions, the forces are initially provided in a sparse format and are not processed by the autoencoder networks.
@@
Decoder: The Decoder takes a flattened tensor of dimension $L$ as its input. Initially, this input tensor undergoes processing via a dense layer. This strategic utilization of the dense layer facilitates a transformation of the arbitrary latent dimensions into a new size that can be conveniently reshaped into a grid structure, a format that seamlessly aligns with the subsequent convolutional layer operations. For example with using a dense layer with 640 units as shown in the Figure 5 allows us to reshape the flat tensor to a $64\times 5\times 2$ tensor, on which series of upsampling and convolution operations are performed until the original spatial dimensions is achieved. In the end, a 2 channel convolution with $1\times 1$ filter is applied to get back the original mesh shape.
Convolutional neural networks are inherently designed to work with grid-like inputs, such as structured meshes. As a result, this methodology isn’t immediately adaptable to inputs that lack this grid-based nature. To address this limitation, we introduce an alternative solution in the form of a fully connected autoencoder network as illustrated in Figure 6. This network can handle diverse and unstructured meshes.
@@ 3.2.2 Gaussian process details
All the Gaussian Process (GP) implementations presented in this work are performed using Matérn covariance functions, as described below:
@@
[…]
The kernel hyperparameters are obtained by training GPs using the scikit-learn library [60] on the latent datasets, as described in Section 2.2. Optimization is performed using the L-BFGS-B optimizer [61], with the number of optimizer restarts chosen as 9 for both cases.
As mentioned earlier, the output of the autoencoder (i.e., the latent representations) is a multioutput function. Therefore, the GP must map the input space $\mathbf{f}$ into a vector function $\mathbf{u}^{l}$. Different methods for such mapping are given in [62]. In this contribution, we model the Gaussian Process Regression (GPR) without considering any correlations between the components of the multidimensional output, implying that we can model the mapping from the inputs to each output component as an independent GP. Note that the outputs of the autoencoders ($\mathbf{u}^{l}$) do not have any physical meaning; hence, we can not make any judgment about their inter-correlation based on physics. However, the results in [37] demonstrate that independent fields can capture the response of multioutput functions with a reasonable accuracy.
3.3 Predictions of the GP+autoencoder framework
We analyse the test example of 2D beam case, with the maximum nodal displacement of 1.16 m. Figure 7a shows the displacement field prediction obtained using the proposed framework which is in great agreement with the FEM solution as presented in the Figure 7b. Figure 7c provides nodal error of the entire framework. While Figure 7d and Figure 7e provide reconstruction error and GP error components as described by Eq.(16) and Eq.(17) respectively. Error plots show that all the errors are fairly low when compared to the magnitude of the displacement solution thereby displaying the capability of the framework in making accurate predictions. Figure 7f provides nodal uncertainty estimates predicted by the framework (2 standard deviations), as obtained from Eq.(14). One can observe that the GP component errors are within the uncertainty bounds for respective nodal predictions, i.e., in other words, true solution of the framework is contained within the displacement distribution for all the nodes.

### v2

@@ Abstract
Many real-world applications demand accurate and fast predictions, as well as reliable uncertainty estimates. However, quantifying uncertainty on high-dimensional predictions is still a severely under-investigated problem, especially when input-output relationships are non-linear. To handle this problem, the present work introduces an innovative approach that combines autoencoder deep neural networks with the probabilistic regression capabilities of Gaussian processes. The autoencoder provides a low-dimensional representation of the solution space, while the Gaussian process is a Bayesian method that provides a probabilistic mapping between the low-dimensional inputs and outputs. We validate the proposed framework for its application to surrogate modeling of non-linear finite element simulations. Our findings highlight that the proposed framework is computationally efficient as well as accurate in predicting non-linear deformations of solid bodies subjected to external forces, all the while providing insightful uncertainty assessments.
@@ 1 Introduction
Rapid advancements in machine learning has naturally fueled developments to approximate modelling. There has been a widespread adoption of machine learning techniques to accelerate computationally intensive numerical methods [11, 12, 13, 14, 15]. In particular, deep learning (DL) models have been successfully used as accurate and computationally inexpensive surrogates to solve non-linear problems in mechanics [16, 17, 18, 19, 20]. In context of ROM, DL methods are being extensively used to find non-linear reduced order representations of the high dimensional data [21, 22]. At the same time, they are computationally inexpensive at the prediction stage, hence, they are highly suitable in a wide area of applications requiring real time simulations [23, 24].
@@
The framework proposed in this contribution focuses on alleviating part of the scaling issue by directly reducing the dimension of outputs, $M$. This is achieved by applying the compressed latent representation of outputs via, so called, autoencoder deep neural networks—a strategy akin to approaches found in related works [44, 45]. Autoencoder neural networks, as an unsupervised learning technique, have been commonly adopted to efficiently learning compressed representations across various domains [46, 47, 48], also in the context of computational fluid dynamics [49, 50] and computational solid mechanics [51, 52]. The primary advantage of autoencoder DNNs is their ability to reduce the dimensionality of data in a non-linear way [53]. As such, they perform much better than POD or PCA when high non-linearities are present in data, see, e.g., the study [54]. Despite the known drawback of DNNs vs. PCA regarding their reduced interpretability of latent representations, in the present work we use DNNs as an integral component of the proposed probabilistic framework. The reason for that is to provide a more general framework, capable of handling a wider class of datasets.
In the present work, we combine GPR with autoencoder DNNs to provide surrogate model for force-displacement predictions in mechanics of solids. In the proposed framework, we reduce the high-dimensional 2D/3D finite element solutions (full-field displacement) into their reduced representation by applying autoencoder neural networks, and then use GPR to provide probabilistic mapping between the input forces and the reduced-order displacements. During the prediction phase, for a given unseen force, GPR predicts the reduced-order (latent space) displacement which is then projected to the full field using the decoder part of autoencoder network. Such general idea to combine GPR with autoencoder DNNs is not new, and has been recently utilized in different applications and contexts. For instance, in the context of high-dimensional outputs, similar strategy has been applied to the spatiotemporal climate modeling [45], in which the full-field climate data is compressed using a DNN autoencoder while GP is used to forecast weather one step ahead. In the context of mechanics, in the works [55, 56], techniques combining GPR and autoencoders have been developed for reliability analysis, and demonstrated for the problem of predicting structural compliance of 2D beams. In the present work, we extend the application to non-linear deformation of 3D solids with unstructured meshes. We emphasize on studying the contribution of errors by individual components (GP and autoencoder), and we provide interpretation of uncertainties in relation to the extrapolated region. We demonstrate the robustness and versatility of our framework by applying it to synthetic datasets generated from the non-linear finite element simulations. Additionally, the source codes and datasets utilized in this work are provided in the following repository: https://github.com/saurabhdeshpande93/gp-auto-regression.
@@ 2.1 Obtaining latent representations with autoencoder neural networks
${\bf u}^{l}=\mathcal{U}_{\text{encoder}}({\bf u},\bm{\theta_{\text{auto}}}^{*}).$ (5)
@@ 2.2 Gaussian Process Regression in the latent space
As discussed in the introduction, Gaussian Process Regression (GPR) faces scalability challenges as the number of data points, $N$, and the dimension of outputs, $M$, increase. In this work we focus on reducing the bottleneck related to $M$, which is particularly important for problems with a large number of degrees of freedom, such as the examples studied in Section 3. As explained earlier, the main idea behind reducing $M$ is to project the high-dimensional displacements (of dimension $M$) to a significantly lower-dimensional space, $L\ll M$, and then employ GPR between the input forces and the latent representation of displacements. In addition to that, in this work we follow a common practice to implement independent GPs for each dimension of the output, assuming that all output components are independent of each other. This approach significantly reduces the $\mathcal{O}(N^{3}L^{3})$ complexity of a multi-output GPR problem to $L$ independent problems of $\mathcal{O}(N^{3})$ complexity each (i.e., $\mathcal{O}(N^{3}L)$). Because our framework uses independent GPRs for each latent output, in this subsection, we provide a brief introduction to single-output GPs. For more details, interested readers are referred to [30].
@@
In Equation (12), $p(\mbox{\boldmath{$y$}}|{\bf F},\bm{\theta_{\text{GP}}},\sigma^{2})$ represents the conditional likelihood of the observations. This likelihood depends on the parameter set $\bm{\theta_{\text{GP}}}$, the input matrix ${\bf F}$, and the noise variance $\sigma^{2}$. The optimised parameters ($\bm{\theta_{\text{GP}}}$, $\sigma^{2}$) are computed by maximising the log likelihood, as discussed in [59]. Note also that the GP and the autoencoder network are trained separately. Autoencoder network needs to be trained first in order to generate the latent state dataset, which is then used to train the GP.
@@ 2.3 Projecting latent GP predictions to the full field space using decoder
As depicted in the Figure 2, the latent space predictions obtained using GP are projected to the full space using decoder part of the autoencoder network. At the inference stage, for a new input force ${\bf f^{*}}$, the GPR first generates latent displacement distribution $p({\tilde{u}}^{*l}|{\bf f^{*}})$ for each latent output, $l$, which is a Gaussian distribution with the mean and variance obtained using Eq.(10) and Eq.(11) respectively:
$p({\tilde{u}}^{*l}|{\bf f^{*}})=\mathcal{N}({\bf\mathbb{E}}_{\text{GP}},{\bf\mathbb{V}}_{\text{GP}}).$ (13)
Subsequently, the distributions of full field displacements are obtained by translating samples from the latent predictive distributions through the decoder component. Let the latent vector ${\bf u}_{s}^{l}$ denote a sample, $s$, drawn from distributions corresponding to all latent outputs, obtained using independent GPs for an input ${\bf f}^{*}$. The displacement corresponding to the full field solution for each degree of freedom (DOF) is a Gaussian distribution, whose mean and variance are computed as follows:
$\begin{split}(\tilde{u}^{*}_{\mu})_{i}&\approx\frac{1}{S}\sum_{s=1}^{S}(\mathcal{U}_{\text{decoder}}({\bf u}_{s}^{l}))_{i},\\ (u_{\sigma}^{*})_{i}^{2}&\approx\frac{1}{S-1}\sum_{s=1}^{S}((\mathcal{U}_{\text{decoder}}({\bf u}_{s}^{l}))_{i}-(u^{*}_{\mu})_{i})^{2},\end{split}$ (14)
@@ 2.4 Finite element formulation for non-linear deformations of solid bodies
Remark: The vector ${\bf f}$ provides a compressed information about the external loading. In our benchmark examples, a body can be subjected to either point loads applied to a single node, or uniform body forces acting throughout the entire volumes. In the 2D example, we apply point loads, and describe the loading ${\bf f}$ as its magnitude ($f_{x}$, $f_{y}$) and the position, $d$, at which the force is applied (relative to the fixed boundary). In the 3D example, we apply body forces, and the loading ${\bf f}$ represents the force density components ($b_{x}$, $b_{y}$, $b_{z}$). See also more details in Section 3.1.
@@ 3 Results
The performance of the framework is demonstrated and studied based on two benchmark examples, as illustrated in Figure 3: the 2D beam example and the 3D liver example. The bodies are subjected to varied loading conditions and undergo large deformations in non-linear regime. The data necessary for training is generated through the finite element simulations, see Section 2.4, and the procedure and data range is described in a more detail in Section3.1. The finite element simulations are performed with the AceFEM framework [60, 61]. The details of the surrogate framework and its training are provided in Section 3.2. In Sections 3.3-3.6 we study the predictive capabilities of the framework, with a more detailed discussion on sources of errors and on the interpretation of probabilistic predictions. In particular, in Sec. 3.6, we demonstrate the capabilities of our framework to be a convenient indicator of the missing data region.
Figure 3: Schematics of examples considered in this work. (a) 2D beam discretised with quad elements is subjected to point loads on the nodes lying on red line; this example is adapted from [57]. (b) 3D liver is subjected to body forces.
@@ 3.1 Dataset generation information
In the 2D beam example, see Figure 3a, a beam is loaded with point forces of random direction and magnitude, within the ranges described in Table 1. These forces are applied to the nodes located along the top line, which is visually represented by red line in the Figure 3a. For the 2D beam case, displacement data is provided for all degrees of freedom (DOFs) of the mesh, capturing the full-field displacements of the beam. However, the forces used as inputs to the surrogate framework are represented in a compressed format of a 3-component array. Specifically, the applied point load is characterized by two components, namely $f_{x}$ and $f_{y}$, representing the force magnitudes in the x and y directions, respectively. Additionally, the distance, $d$, from the fixed boundary to the point of application is considered. This information is visually depicted in Figure 3a.
In the 3D liver example, a body representing the liver geometry is subjected to a body force of random direction and magnitude. The liver is also constrained at specific nodes located at the right end, see Figure 3b. Similarly to the beam case, displacement data is provided for all DOFs of the liver mesh, representing the full-field displacements of the organ. The compressed representation of forces consists of the body force density components in the x, y, and z directions, respectively ($b_{x},b_{y},b_{z}$). Further details, such as the range of external forces applied, as well as other information used to generate the datasets are given in the Table 1.
Remark: Note that for each loading case, the data generated by the FE simulation is deterministic. This has been a choice in the scope of the present work, however, it can be extended to the non-deterministic case if needed.
@@ 3.2.1 Autoencoder networks: Architectures & training details
In this study, we employ two types of autoencoder networks. Firstly, we utilize a CNN-based architecture, illustrated in Figure 4, tailored for the 2D beam example. Secondly, we utilize a fully connected autoencoder network, depicted in Figure 5, implemented for the 3D liver example. Both of these networks consist of an encoder and a decoder component. The primary function of these autoencoder networks is to compress the displacement solutions of the full field space into corresponding latent representations. Note that the forces are initially provided in a simple format, see Sec. 3.1, thus do not need to be compressed by autoencoder networks.
@@
Decoder: The Decoder takes a flattened tensor of dimension $L$ as its input. Initially, this input tensor undergoes processing via a dense layer. This strategic utilization of the dense layer facilitates a transformation of the arbitrary latent dimensions into a new size that can be conveniently reshaped into a grid structure, a format that seamlessly aligns with the subsequent convolutional layer operations. For instance, utilizing a dense layer of 640 units as shown in the Figure 4 allows us to reshape the flat tensor to a $64\times 5\times 2$ tensor, on which series of upsampling and convolution operations are performed until the original spatial dimensions is achieved. In the end, a 2 channel convolution with $1\times 1$ filter is applied to get back the original mesh shape.
@@
Convolutional neural networks are inherently designed to work with grid-like inputs, such as structured meshes. For that reason, this methodology isn’t immediately adaptable to inputs that lack this grid-based nature, such as the unstructured mesh that are present in the liver example. To address this limitation, we introduce an alternative solution in the form of a fully connected autoencoder network as illustrated in Figure 5. This network can handle diverse and unstructured meshes.
@@
Remark: Besides the size of the latent vector, there are several other hyperparameters that determine performance of both autoencoder architectures. In the case of 2D CNN network, these are: the size of convolution window, the number of channels per layer, the number of convolutions per level, and the type of pooling layers. In the case of 3D fully connected network, these are: the sizes of each layer, and the arrangement of skip connections. The particular choice of hyperparameters that determine both autoencoder architectures is primarily based on our earlier experience and studies, see [57, 19], which is then followed by the grid search optimization in the space of hyperprameters. The presented final architectures have been found to perform the best for the problems analyzed in this work.
@@ 3.2.2 Gaussian process details
All the Gaussian Process (GP) implementations presented in this work are performed using the Matérn covariance function, as it is known to be less prone to overfitting. (Note, however, that our framework is not limited to any particular covariance function.) The Matérn covariance function has the following form:
$C_{\nu}(r_{ij})=\lambda^{2}\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\sqrt{2\nu}\frac{r_{ij}}{\psi}\right)^{\nu}K_{\nu}\!\left(\sqrt{2\nu}\frac{r_{ij}}{\psi}\right),$ (18)
@@
$k(r_{ij})=C_{5/2}(r_{ij})=\lambda^{2}\Big{(}1+\frac{\sqrt{5}r_{ij}}{\psi}+\frac{5r_{ij}^{2}}{3\psi^{2}}\Big{)}\text{exp}\bigg{(}-\frac{\sqrt{5}r_{ij}}{\psi}\bigg{)}.$ (19)
The remaining hyperparameters, $\psi$ and $\lambda^{2}$, are then inferred or optimized by maximizing the marginal likelihood in Eq. (12).
The kernel hyperparameters are obtained by training GPs using the scikit-learn library [65] on the latent datasets, as described in Section 2.2. Optimization is performed using the L-BFGS-B optimizer [66]. As described in Section 2.2, we implement independent GPs corresponding to each latent component of the autoencoder network.
3.3 Sources of prediction errors
At the prediction phase, for a new unseen case, the GPR is first applied to achieve the latent space prediction, which is subsequently projected to the full space solution using the decoder part of the autoencoder network, see Section 2 and Figure 2. Hence the error in the framework arises from GPR fitting as well as the compression and reconstructions steps done by the autoencoder network. The idea of this split of errors is schematically shown in Figure 6 for a single DOF of the full-field solution.
The mean prediction of the framework, $\tilde{u}_{\mu}$, is interpreted as the displacement solution. Hence the absolute difference between this mean prediction, $\tilde{u}_{\mu}$, and the true FEM solution, $u$, is considered as the error of the framework,
$e_{\text{f}}=|\tilde{u}_{\mu}-u|.$ (20)
The error contributed to autoencoder, $e_{r}$, is computed by taking absolute difference between the reconstructed solution and the FEM solution. The reconstructed solution is obtained by projecting the true latent solution (as obtained by Eq. (5)) to the full field, thus neglecting any contribution of the GP. The reconstructed solution reads ${\bf u}_{r}=\mathcal{U}_{\text{decoder}}({\bf u}^{l})$, and the corresponding reconstruction error is
$e_{r}=|u_{r}-u|.$ (21)
Finally, the error contributed to GP fitting, $e_{\text{GP}}$, is computed by taking the difference between the mean full-field prediction, $\tilde{u}_{\mu}$, and the reconstruction solution, $u_{r}$, as follows:
$e_{\text{GP}}=|\tilde{u}_{\mu}-u_{r}|.$ (22)
Note: All the errors are computed in the decoded space, and are provided per full-field DOF.
3.4 Performance study of the GP+autoencoder framework
In this section we study the performance of the framework. Firstly, we showcase two particular load cases for the 2D and 3D example representing highest nodal displacement examples of each scenario. We visualize the predictions and corresponding errors, which are plotted on deformed meshes. In particular, we utilize the aforementioned breakdown of the framework error into reconstruction- and GP error, which aims to provide a clear and accessible understanding of how each component contributes to the overall performance of the framework. Figures 7 and 8 illustrate the interpolated node-wise L2 norm of the framework’s predictions, including displacements, associated uncertainties, and errors for the 2D beam and 3D liver cases, respectively. Further, we provide overall error metrics for the entire test sets to quantify the framework’s performance.